Lab Question: Reflection & Refraction (Snell's Law) experiment

1) If you conducted this experiment with a different lens, one with larger refractive index, would you find larger angles of refraction for each angle incidence? Why or why not??


2) Without refrence to equations, explain why it's physically impossible for total internal reflection (theta= 90 degree) to occur in this experiment ????
Answer #1

I do not see any image. Just guessing that refracted light travels through the medium of lens and not the incident light.

(1)

According to Snell's law,

n_{1} * sin\theta _{1} = n_{2} * sin\theta _{2}

Now, suppose we conduct the same experiment with different lens with larger refractive index n2. The value of n1 remains the same. Hence, for the Snell's law to hold good, value of sin\theta _{2} should decrease with increasing value of angle of incidence. Since the value of sin\theta _{2} decrease with the decreasing value of \theta _{2}, no, we would not find larger angles of refraction for each angle of incidence. Instead, with the increasing value of incidence, the angles of refraction would decrease.

(2)

For the total internal reflection to happen, the light wave should be traveling in a medium with higher refractive index and with an incident angle greater than the critical angle for the medium and the medium on the other side of the boundary . Since this is not the case in this experiment, the total internal reflection will not happen.